Phyllis invested $10,000, a portion earning a simple interest rate of 3 1/5%
per year and the rest earning a rate of 3% per year. After 1 year the total interest earned on these investments was $312. How much money did she invest at each rate?

Answer by mananth(16949)

Investment Part I 3.50% per annum ----x
Investment part II 3.00% per annum ----y
The sum of the investments is $10,000.00
The sum of individual interests = $312.00
x+y= 10000 ------------------------1
3.50%x+ 3.00%y =$312.00
Multiply by 100
3.5x+3y =$31,200.00 --------2

Multiply (1) by -3.5
we get

-3.5x -3.5y= -35000.00

Add this to (2)

0x-0.5y =-$3,800.00

divide by-0.5

y = $7,600.00 investment at 3.00%
Balance $2,400.00 investment at 3.50%
CHECK
$2,400.00 @ 3.50% $84.00
$7,600.00 @ 3.00% $228.00
Total -------------------- $312.00